Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 78
Entry time: 12 Aug 2019, 09:50
|
Author: |
Gerd Wellenreuther |
Location: |
General |
Severity: |
Information |
Subject: |
Status report after dispatch meeting 12.8. |
|
|
XO
FXE
- XTD1 CRLs: There is an issue regarding a limit switch, to be repaired in the shutdown by MKnoll + EEE.
SPB/SFX
- Master loop has to be updated. EEE to agree offline with SPB/SFX ==> has been fixed
- Hexapods are being integrated. ==> was supposed to be tackled today. But since both HED and MID experienced different issues with their hexapods, the commissioning at SPB/SFX has to rescheduled (Tuesday?)
- Micro KB encoders + beam are drifting away, so closed loop operation would stabilize things... has to be commissioned, foreseen for Wednesday. Users start at Thursday...
- A microscope got stuck on a limit switch. Probably the limit switches have to be rewired.
- An encoder in loop 11 seems to be missing. EEE is on it...
SCS
- Since Sunday there appears to be an issue with the in-vacuum sample stage. Is kind of urgent...
SQS
- Last week: Commisioned the laser
- This week: Users & night shift
MID
- There is an issue with the hexapod, EEE is on it.
- Chris is testing some cooling / heating using peltier elements.
HED
- Train IDs are missing in loop 5 and 10. EEE is on it, too.
Optical lasers
VAC
XRO
XPD
- There is probably an intervention in XTD1 to repair a camera for the self-seeding setup. Others might want to jump on the train. Time is not yet decided.
DET
EEE
CAS
- Test for 2.6. are ongoing.
ITDM
SRP
TS
AoB
|